A.CH3(CH2)14COO

B.CH3(CH2)5CH=CH(CH2)7COO

How Many ATP will be formed after beta- oxidation of A and B?

i dont think you need them. i just want what to know the ATP difference between saturated and unsaturated compound

16 c(palmitoleic (unsaturated) , 16c palmitic acid (saturated)
Answer #1

Complete oxidation of one molecule of palmitic acid yields 129 ATP molecules

C15H31COOH+8CoASH+ATP+7FAD+7NAD++7H2O → 8CH3COSCoA+AMP+PPi+7FADH2+7NADH+7H+

Each molecule of acetyl-CoA yields 12 ATP (12 × 8 = 96); FADH2 yields 2 ATP (7 × 2 = 14); NADH yields 3ATP (7 × 3 = 21); and two high-energy bonds are consumed (-2; ATP → AMP + PPi). Thus, net ATP production is 129.

Beta-oxidation of unsaturated fatty acids changes the ATP yield due to the requirement of two possible additional enzymes. Beta oxidation of unsaturated fatty acid skips one FAD → FADH2 step hence has lower ATP production than saturated fatty acid.